A market has a linear demand schedule with a slope of −0.3. When price is £3, quantity sold is 30 units. Where does this demand schedule hit the price and quantity axes? What is price if quantity sold is 25 units? How much would be sold

at a price of £9?


1
Expert's answer
2021-01-21T18:15:05-0500

a) The equation of the demand function is


"P=12-0.3Q,"

here, -0.3 is the slope of the demand function.

When Q=0 then P=£12. If P=0 then Q=40. Therefore, this demand schedule hit the price and quantity axes at P=12 and Q=40, respectively.

b) Let's calculate the price if quantity sold is 25 units:


"P=12-0.3\cdot25=£4.5"

c) Let's calculate the quantity of units sold at a price of £9:


"9=12-0.3Q,""Q=\dfrac{12-9}{0.3}=10"